# Shadows and Observational Images of a Schwarzschild-like Black Hole Surrounded by a Dehnen-type Dark Matter Halo

### v1

@@ III.1 The trajectory of photon motion and the effective potential
Therefore, in Fig. 4, we present the constrained parameter space of $\rho_{s}$ and $r_{s}$ for a Schwarzschild-like BH surrounded by a Dehnen-type DM halo, based on the limits of the shadow radius of M87*. The black dashed line in Fig. 4  represents the BH shadow radius $r_{sh}=6.25$. Note that all values in the density plot are greater than the lower bound of the shadow radius, i.e., $r_{sh}>4.75$. The region enclosed by this line and the $\rho_{s}$ and $r_{s}$ axes corresponds to shadow radii smaller than 6.25, indicating the allowed range of $\rho_{s}$ and $r_{s}$. Hence, the parameter values of $\rho_{s}$ and $r_{s}$ in the following analysis all lie within this range.
Figure 4: The constraint imposed by the shadow radius of the M87* on the DM halo density $\rho_{s}$ and radius $r_{s}$. The dashed line represents the upper limit of the shadow radius, which is 6.25.
IV The images of a Schwarzschild-like BH with static thin accretion disks.
In this section, we aim to study the optical appearance of a Schwarzschild-like BH surrounded by a Dehnen-type DM halo and illuminated by a static geometrically thin accretion disk. We assume that the observer is located at the North Pole, directly facing the accretion disk on the equatorial plane. Since photons emitted from different regions of the disk contribute differently to the BH image, we first classify the photon trajectories around the Schwarzschild-like BH and then analyze the resulting image.
IV.1 Classification of photon trajectories: direct emission, lensing ring, photon ring.
Photons emitted from the accretion disk that reach a distant observer can equivalently be regarded as light rays traced backward from the observer that intersect the accretion disk near the BH. Each time a photon intersects with the accretion disk, it will extract energy from the accretion disk. Based on this, we adopt the classification method proposed in the literature Gralla et al. (2019) and divide the photon trajectories into three categories based on the number of intersections $n$ with the accretion disk.
@@
here $\phi$ denotes the total change in the azimuthal angle over the course of the entire photon trajectory. From Eq. (16), it can be known that the number of intersections $n$ is actually a function of the impact parameter $b$, and there is a specific relationship that they satisfy Peng et al. (2021):
@@
For a given value of the $m$, there exist $b_{m}^{\pm}$. Here $b_{m}^{-}<b_{c}$ and $b_{m}^{+}>b_{c}$. Hence, null geodesics can be classified into three categories: direct emission, the lensing ring, and the photon ring. And they can be expressed as follows:
  1. (1)
    Direct emission: 1/$41/4<n<3/$ 4 $\Leftrightarrow$ $b\in(0,b_{2}^{-})\cup(b_{2}^{+},\infty)$;
  2. (2)
    Lensing ring: 3/$43/4<n<5/$ 4 $\Leftrightarrow$ $b\in(b_{2}^{-},b_{3}^{-})\cup(b_{3}^{+},b_{2}^{+})$;
  3. (3)
    Photon ring: $n>5/4$Leftrightarrow $b\in(b_{3}^{-},b_{3}^{+})$.
@@
Furthermore, in the left panel of Fig. 6, we show how the intersection number $n$ varies with the impact parameter $b$ for different values of $r_{s}$ with fixed $\rho_{s}=0.01$, while in the right panel, we fix $r_{s}=0.06$ and show the variation of $n$ with $b$ for different values of $\rho_{s}$. It can be found that when the value of $\rho_{s}$ is fixed, as the parameter $r_{s}$ increases, the corresponding curves as a whole show a trend of shifting to the right. This phenomenon illustrates that as the DM halo radius $r_{s}$ increases, the range of the impact parameter $b$ associated with the lensing and photon rings also expands. Similarly, when $r_{s}$ takes a fixed value, the impact parameter of the system also shows a variation rule of increasing monotonically with the increase of $\rho_{s}$.
@@
Moreover, through further calculations, we have given the impact parameter intervals with the number of intersections with the accretion disk. These interval values are presented in Tables 1 and 2. The table also provides the calculations for various parameters, including the radius of the innermost stable circular orbit $r_{isco}$, the photon sphere radius $r_{ph}$, the event horizon radius $r_{h}$, and the critical impact parameter $b_{ph}$The radius of the innermost stable circular orbit can be determined using the following equation. Wang et al. (2023)
$\begin{split}r_{isco}=\frac{3f(r_{isco})f^{\prime}(r_{isco})}{2(f^{\prime}(r_{isco}))^{2}-f(r_{isco})f^{\prime\prime}(r_{isco})}\end{split}$ (19)
@@
It is clear that in Tables  1 and  2, all relevant physical quantities increase with the DM parameters.
IV.2 Transfer functions
@@
We denote the transfer function as $r_{m}(b)$ ($m=1,2,3,\cdots$), which represents the location of the $m$-th intersection point between the light ray and the accretion disk. The first three transfer functions can be formulated as what is shown below Tsukamoto (2022)
@@
Since the observer measures the average brightness proportional to the flux Gralla et al. (2019), the slope of the transfer function $\frac{dr_{m}}{db}$ represents the demagnification factor of the optical image. A larger slope corresponds to a narrower interval of the impact parameter, resulting in a smaller contribution to the total observed intensity of the BH. Theoretically, photons in the photon ring region can intersect the accretion disk many times, but due to the steep slope of their transfer functions, their contribution to the observed intensity is almost negligible. Therefore, we only consider the first three transfer functions here.
IV.3 Optical appearance of Schwarzschild-like black hole with thin accretion disk.
As discussed in the previous subsection, we consider an optically thin and geometrically thin, static accretion disk in the equatorial plane of a Schwarzschild-like BH as the sole radiation source, with isotropic emission. Therefore, the specific intensity of radiation emitted at frequency $\nu_{e}$, as measured by a distant observer at the north pole, is given by Wang et al. (2023):
@@
[…]
@@
Here, $r_{m}(b)$ is the transfer function. To observe the shadow and optical appearance of a Schwarzschild-like BH surrounded by a Dehnen-type DM halo, we consider projecting the optical appearance onto a two-dimensional plane under different emission models. Since photon motion along the unstable photon sphere constitutes the boundary of the BH shadow, the shadow shape can be obtained by projecting it from the BH plane to the observer’s plane with celestial coordinates $(x,y)$. Here we define the celestial coordinate system as Kumar and Ghosh (2020)
@@
Therefore, it suffices for us to define the emission intensity of the accretion disk and calculate the corresponding received intensity using Eq. 25, and then project this intensity onto the two-dimensional plane in the celestial coordinate system to obtain the optical appearance of the BH. We consider three static accretion disk models:
@@
    here, $I_{0}$ is the maximum emission intensity, and $r_{isco}$ is the radius of the innermost stable circular orbit defined previously.
@@
    $\begin{split}I_{em}(r)=\begin{cases}I_{0}\frac{\frac{\pi}{2}-\arctan\theta\left[r-(r_{isco}-1)\right]}{\frac{\pi}{2}-\arctan\theta\left[r_{h}-(r_{isco}-1)\right]},&r>r_{h},\\ 0,&r\leq r_{h}.\end{cases}\end{split}$ (30)
@@
Figure 9: Three types of emission intensity (left column), received intensity (middle column), and optical appearance (right column) of the Schwarzschild-like BH under different DM halo parameter values. We consider the case of fixed $\rho_{s}=0.01$ with $r_{s}=0.1$, 0.15, and 0.2. The images in the right column are composed of the optical appearances for the Schwarzschild BH (top left) and the Schwarzschild-like BH with $r_{s}=0.1$ (top right), $r_{s}=0.15$ (bottom right), and $r_{s}=0.2$ (bottom left).
@@
  • As the DM halo radius $r_{s}$ gradually increases, both the emission intensity and the received intensity exhibit a rightward shift. This phenomenon suggests that several important physical quantities of the Schwarzschild-like black hole surrounded by a Dehnen-type DM halo — including the radius of the innermost stable circular orbit , the photon sphere radius, and the event horizon radius — also increase progressively with increasing $r_{s}$.
@@
Figure 10: Three types of emission intensity (left column), received intensity (middle column), and optical appearance (right column) of the Schwarzschild-like BH under different DM halo parameter values. We consider the case of fixed $r_{s}=0.06$ with $\rho_{s}=0.1$, 0.15, and 0.2. The images in the right column are composed of the optical appearances for the Schwarzschild BH (top left) and the Schwarzschild-like BH with $\rho_{s}=0.1$ (top right), $\rho_{s}=0.15$ (bottom right), and $\rho_{s}=0.2$ (bottom left).
V The images of a Schwarzschild-like BH with static thin spherical accretions
In previous section, the thin accretion disk considered in this study represents a simplified theoretical model. It is based on the assumption that after matter existing in the universe gets captured by a black hole, it gradually forms the shape of a thin accretion disk. However, when the accreting matter possesses relatively low angular momentum, the scenario differs fundamentally. In this case, instead of forming a thin accretion disk, the matter will accrete radially onto the BH, ultimately establishing a spherical accretion state. Building upon this framework, we investigate the optical appearance of a Schwarzschild-like BH embedded in a Dehnen-type DM halo under spherical accretion. Specifically, we examine how variations in the DM halo’s density $\rho_{s}$ and radius $r_{s}$ influence the observational features.
The radiation intensity observed by an observer at infinity, when a BH is surrounded by spherical accretion, can be calculated by integrating the specific intensity of photon radiation along the path $\gamma$ Bambi (2013); Jaroszynski and Kurpiewski (1997)
[…]
@@ V The images of a Schwarzschild-like BH with static thin spherical accretions
[…]
here, $\frac{d\phi}{dr}$ can be obtained from the equation of motion (Eq. (12)) of the photon in the gravitational field. Then, substitute it into Eq. (34).
[…]
@@
Figure 11: The graph depicting how the total observed intensity changes with the impact parameter in the context of the spherical accretion model. On the left side are the images when the DM halo density $\rho_{s}=0.01$ and $r_{s}$ takes the values of 0.1, 0.15, 0.2, and Schwarzschild case. On the right side are the images when $r_{s}=0.06$ and $\rho_{s}$ takes the values of 0.1, 0.15, and 0.2, and Schwarzschild case.
As shown clearly in Fig. 11, when one of the DM halo parameters (either $\rho_{s}=0.01$ or $r_{s}=0.06$) is fixed, the observed intensity steadily diminishes as the other parameter increases, and the overall curve shifts to the right. The peak value of the observed intensity appears at the position where $b=b_{ph}$, and as the variable gradually increases, the position where the peak value appears also shows a trend of gradually increasing. At the peak position, photons orbit multiple times to generate the maximum intensity. As the parameter increases, the observed intensity gradually diminishes, indicating that the brightness of the BH image decreases correspondingly with increasing parameter values. Furthermore, the BH image illuminated by spherical accretion is presented in Fig. 12.
@@
VI conclusion

### v2

@@ 3.2 Constraints of the M87* shadow on parameters
Therefore, in Fig. 4, we present the constrained parameter space of $\rho_{s}$ and $r_{s}$ for a Schwarzschild-like BH surrounded by a Dehnen-type DM halo, based on the limits of the shadow radius of M87*. The black dashed line in Fig. 4  represents the BH shadow radius $r_{sh}=6.25$. It should be noted that all values in the density plot exceed the Schwarzschild BH’s shadow radius (i.e., $r_{sh}>3\sqrt{3}$). Moreover, the region enclosed by this line and the $\rho_{s}$ and $r_{s}$ axes corresponds to shadow radii smaller than 6.25, indicating the allowed range of $\rho_{s}$ and $r_{s}$. Hence, the parameter values of $\rho_{s}$ and $r_{s}$ in the following analysis all lie within this range.
Figure 4: The constraints imposed by the M87* shadow radius on the DM halo density $\rho_{s}$ and radius $r_{s}$. The dashed line indicates the upper limit of the shadow radius (6.25), with the parameter space below this line representing the allowed region for $\rho_{s}$ and $r_{s}$.
4 The images of a Schwarzschild-like black hole with static thin accretion disks
In this section, we aim to study the optical appearance of a Schwarzschild-like BH surrounded by a Dehnen-type DM halo and illuminated by a static geometrically thin accretion disk. We assume that the observer is located at the North Pole, directly facing the accretion disk on the equatorial plane. Since photons emitted from different regions of the disk contribute differently to the BH image, we first classify the photon trajectories around the Schwarzschild-like BH and then analyze the resulting image. The trajectories of photons have been partially investigated in Rani:2025esb . Therefore, within the parameter ranges constrained by observational data, we focus here on briefly analyzing the influence of DM parameters on photon trajectories. Additionally, we employ three distinct thin disk accretion models to study the corresponding image features under their illumination.
4.1 Classification of photon trajectories: direct emission, lensing ring, photon ring
Photons emitted from the accretion disk that reach a distant observer can equivalently be regarded as light rays traced backward from the observer that intersect the accretion disk near the BH. Each time a photon intersects with the accretion disk, it will extract energy from the accretion disk. Based on this, we adopt the classification method proposed in the literature Gralla:2019xty and divide the photon trajectories into three categories based on the number of intersections $n$ with the accretion disk.
@@ 4.1 Classification of photon trajectories: direct emission, lensing ring, photon ring
here $\phi$ denotes the total change in the azimuthal angle over the course of the entire photon trajectory. From Eq. (17), it can be known that the number of intersections $n$ is actually a function of the impact parameter $b$, and there is a specific relationship that they satisfy Peng:2020wun :
@@
For a given value of the $m$, there exist $b_{m}^{\pm}$. Here $b_{m}^{-}<b_{ph}$ and $b_{m}^{+}>b_{ph}$. Hence, null geodesics can be classified into three categories: direct emission, the lensing ring, and the photon ring. And they can be expressed as follows:
  • Direct emission: $1/4<n<3/4$ $\Leftrightarrow$ $b\in(0,b_{2}^{-})\cup(b_{2}^{+},\infty)$;
  • Lensing ring: $3/4<n<5/4$ $\Leftrightarrow$ $b\in(b_{2}^{-},b_{3}^{-})\cup(b_{3}^{+},b_{2}^{+})$;
  • Photon ring: $n>5/4$ $\Leftrightarrow$ $b\in(b_{3}^{-},b_{3}^{+})$.
@@
Furthermore, in the left panel of Fig. 6, we show how the intersection number $n$ varies with the impact parameter $b$ for different values of $r_{s}$ with fixed $\rho_{s}=0.01$, while in the right panel, we fix $r_{s}=0.06$ and show the variation of $n$ with $b$ for different values of $\rho_{s}$. It can be found that when the value of $\rho_{s}$ is fixed, as the parameter $r_{s}$ increases, the corresponding curves as a whole show a trend of shifting to the right. This phenomenon illustrates that as the DM halo radius $r_{s}$ increases, the impact parameter $b$ associated with the lensing and photon rings also increases. Similarly, when $r_{s}$ takes a fixed value, the impact parameter of the system also shows a variation rule of increasing monotonically with the increase of $\rho_{s}$.
@@
Moreover, through further calculations, we have given the impact parameter intervals with the number of intersections with the accretion disk. These interval values are presented in Tables 1 and 2. The table also provides the calculations for various parameters, including the radius of the innermost stable circular orbit $r_{isco}$, the photon sphere radius $r_{ph}$, the event horizon radius $r_{h}$, and the critical impact parameter $b_{ph}$ .The radius of the innermost stable circular orbit can be determined using the following equation Wang:2023vcv
$\begin{split}r_{isco}=\frac{3f(r_{isco})f^{\prime}(r_{isco})}{2(f^{\prime}(r_{isco}))^{2}-f(r_{isco})f^{\prime\prime}(r_{isco})},\end{split}$ (20)
where the prime denotes differentiation with respect to the radial coordinate $r$. It is clear that in Tables  1 and  2, all relevant physical quantities increase with the DM parameters. This is consistent with our previous discussion.
@@
4.2 Transfer functions
@@
We denote the transfer function as $r_{m}(b)$ ($m=1,2,3,\cdots$), which represents the location of the $m$-th intersection point between the light ray and the accretion disk. The first three transfer functions can be formulated as what is shown below Tsukamoto:2021apr
@@
Since the observer measures the average brightness proportional to the flux Gralla:2019xty , the slope of the transfer function $\frac{dr_{m}}{db}$ represents the demagnification factor of the optical image. A larger slope corresponds to a narrower interval of the impact parameter, resulting in a smaller contribution to the total observed intensity of the BH. Theoretically, photons in the photon ring region can intersect the accretion disk many times, but due to the steep slope of their transfer functions, their contribution to the observed intensity is almost negligible. Therefore, we only consider the first three transfer functions here.
4.3 Optical appearance of Schwarzschild-like black hole with thin accretion disk
As discussed in the previous subsection, we consider an optically thin and geometrically thin, static accretion disk in the equatorial plane of a Schwarzschild-like BH as the sole radiation source, with isotropic emission. Therefore, the specific intensity of radiation emitted at frequency $\nu_{e}$, as measured by a distant observer at the north pole, is given by Wang:2023vcv :
@@ 4.3 Optical appearance of Schwarzschild-like black hole with thin accretion disk
$\begin{split}I_{obs}(r)=\int I_{0}(r,\nu_{0})d\nu_{0}=\int g^{4}I_{e}(r,\nu_{e})d\nu_{e}=f(r)^{2}I_{em}(r)\end{split}.$ (25)
Among them, we denote $I_{em}(r)=\int I_{e}(r,\nu_{e})d\nu_{e}$ as the total emission intensity. From the above discussion, When photon trajectories are traced backward from the observer, they may intersect the accretion disk multiple times, as can be observed. Each intersection contributes additional intensity from the disk, therefore the total observed intensity at the observer is the sum of intensities from all these intersection points Peng:2020wun ; Gralla:2019xty
@@
Here, $r_{m}(b)$ is the transfer function. To observe the shadow and optical appearance of a Schwarzschild-like BH surrounded by a Dehnen-type DM halo, we consider projecting the optical appearance onto a two-dimensional plane under different emission models. Since photon motion along the unstable photon sphere constitutes the boundary of the BH shadow, the shadow shape can be obtained by projecting it from the BH plane to the observer’s plane with celestial coordinates $(x,y)$. Here we define the celestial coordinate system as Kumar:2018ple
@@
Therefore, it suffices for us to define the emission intensity of the accretion disk and calculate the corresponding received intensity using Eq. (26), and then project this intensity onto the two-dimensional plane in the celestial coordinate system to obtain the optical appearance of the BH. We consider three static accretion disk models:
@@
    where $I_{0}$ is the maximum emission intensity, and $r_{isco}$ is the radius of the innermost stable circular orbit defined previously.
@@
    $\begin{split}I_{em}(r)=\begin{cases}I_{0}\frac{\frac{\pi}{2}-\arctan\left[r-(r_{isco}-1)\right]}{\frac{\pi}{2}-\arctan\left[r_{h}-(r_{isco}-1)\right]},&r>r_{h},\\ 0,&r\leq r_{h}.\end{cases}\end{split}$ (31)
@@
Figure 9: Emission intensity (left column), received intensity (middle column), and corresponding optical appearance (right column) for three models with fixed $\rho_{s}=0.01$. In the left two columns, the blue, orange, green, and red curves represent the Schwarzschild case and the DM parameters with $r_{s}=0.1$, 0.15, and 0.2, respectively. The right panel composites the optical appearances of: Schwarzschild BH (top left), the Schwarzschild-like BH with $r_{s}=0.1$ (top right), $r_{s}=0.15$ (bottom right), and $r_{s}=0.2$ (bottom left).
@@
  • As the DM halo radius $r_{s}$ gradually increases, both the emission intensity and the received intensity exhibit a rightward shift. This phenomenon suggests that several key physical quantities of the Schwarzschild-like BH surrounded by a Dehnen-type DM halo, such as the radius of the innermost stable circular orbit, the photon sphere radius, and the event horizon radius, also increase progressively with increasing $r_{s}$.
@@
Figure 10: Emission intensity (left column), received intensity (middle column), and corresponding optical appearance (right column) for three models with fixed $r_{s}=0.06$. In the left two columns, the blue, orange, green, and red curves represent the Schwarzschild case and the DM parameters with $\rho_{s}=0.1$, 0.15, and 0.2, respectively. The right panel composites the optical appearances of: Schwarzschild BH (top left), the Schwarzschild-like BH with $\rho_{s}=0.1$ (top right), $\rho_{s}=0.15$ (bottom right), and $\rho_{s}=0.2$ (bottom left).
5 The images of a Schwarzschild-like black hole with thin spherical accretions
In previous section, the thin accretion disk considered in this study represents a simplified theoretical model. It is based on the assumption that after matter existing in the universe gets captured by a BH, it gradually forms the shape of a thin accretion disk. However, when the accreting matter possesses relatively low angular momentum, the scenario differs fundamentally. In this case, instead of forming a thin accretion disk, the matter will accrete radially onto the BH, ultimately establishing a spherical accretion state Yuan:2014gma . Building upon this framework, we investigate the optical appearance of a Schwarzschild-like BH embedded in a Dehnen-type DM halo under spherical accretion. Specifically, we examine how variations in the DM halo’s density $\rho_{s}$ and radius $r_{s}$ influence the observational features.
5.1 Optical appearance of Schwarzschild-like black hole with static thin spherical accretions
We first consider a simplified scenario of thin spherical accretion, where the accreting matter is assumed to be statically distributed around this Schwarzschild-like BH. The radiation intensity observed by an observer at infinity, when a BH is surrounded by thin spherical accretion, can be calculated by integrating the specific intensity of photon radiation along the path $\gamma$ Bambi:2013nla ; Jaroszynski:1997bw
$\begin{split}I(\nu_{0})=\int_{\gamma}g^{3}j_{e}(\nu_{e})dl_{prop}\end{split},$ (32)
@@ 5.1 Optical appearance of Schwarzschild-like black hole with static thin spherical accretions
$\begin{split}I_{obs}=\int_{\gamma}\frac{f(r)^{2}}{r^{2}}\sqrt{\frac{1}{f(r)}+r^{2}\left(\frac{d\phi}{dr}\right)^{2}}dr\end{split},$ (35)
here, $\frac{d\phi}{dr}$ can be obtained from the equation of motion (Eq. (12)) of the photon in the gravitational field. Then, substitute it into Eq. (35), we have
$\begin{split}I_{obs}=\int_{\gamma}\frac{g^{4}}{r^{2}}\sqrt{\frac{1}{f(r)}+\frac{b^{2}}{r^{2}-b^{2}f(r)}}dr\end{split}.$ (36)
@@
Figure 11: The total observed intensity changes with the impact parameter in the context of the static spherical accretion model. On the left side are the observed intensities when the DM halo density $\rho_{s}=0.01$ and $r_{s}$ takes the values of 0.1 (blue curve), 0.15 (green curve), 0.2 (black curve), and Schwarzschild (red curve) case. On the right side are the observed intensities when $r_{s}=0.06$ and $\rho_{s}$ takes the values of 0.1 (blue curve), 0.15 (green curve), 0.2 (black curv), and Schwarzschild (red curve) case.
In Fig. 11, when one of the DM halo parameters (either $\rho_{s}=0.01$ or $r_{s}=0.06$) is fixed, the observed intensity steadily diminishes as the other parameter increases, and the overall curve shifts to the right. The peak value of the observed intensity appears at the position where $b=b_{ph}$, and as the variable gradually increases, the position where the peak value appears also shows a trend of gradually increasing. At the peak position, photons orbit multiple times to generate the maximum intensity. As the parameter increases, the observed intensity gradually diminishes, indicating that the brightness of the BH image decreases correspondingly with increasing parameter values. Furthermore, the BH image illuminated by spherical accretion is presented in Fig. 12.
@@
5.2 Optical appearance of Schwarzschild-like black hole with radially infalling spherical accretions
Next, we consider a more general scenario where the Schwarzschild-like BH is surrounded by radially infalling spherical accretion. We adopt the same emission profile as in the static spherical accretion case, which implies that Eq. (33) remains applicable. For the infalling accretion, the redshift factor can be expressed as Bambi:2013nla ; Wang:2023vcv ; KumarWalia:2024yxn :
$\begin{split}g=\frac{P_{\mu}u_{o}^{\mu}}{P_{\nu}u_{e}^{\nu}}.\end{split}$ (37)
Here, $P_{\mu}$ denotes the 4-momentum of photons emitted by the accreting matter, while $u_{o}^{\mu}$ and $u_{e}^{\mu}$ represent the 4-velocities of the observer and emitting matter respectively. For a distant observer, the 4-velocity is given by $u_{o}^{\mu}=(1,0,0,0)$, while for radially infalling accreting matter, the 4-velocity can be expressed as
$u_{e}^{t}=\frac{1}{f(r)},\qquad u_{e}^{r}=-\sqrt{1-f(r)},\qquad u_{e}^{\theta}=u_{e}^{\phi}=0.$ (38)
The photon 4-momentum $P_{\mu}$ can be derived from the Lagrangian via $P_{\mu}=\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}$, yielding
$\begin{split}P_{t}&=-\frac{1}{b},\\ P_{r}&=\pm\frac{1}{f(r)}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}},\end{split}$ (39)
where the $\pm$ sign indicates the photon moving inward or outward relative to the Schwarzschild-like BH. Note that here we only display the two required components of the photon 4-momentum. By substituting Eqs. (38) and (39) into Eq. (37), we can therefore calculate the corresponding redshift factor
$g=\frac{1}{\frac{1}{f(r)}\pm\sqrt{\frac{1}{f(r)}\left[\frac{1}{f(r)}-\frac{b^{2}}{r^{2}}\right][1-f(r)]}}.$ (40)
Regarding the proper length $dl_{prop}$, it is defined as Bambi:2013nla ; Wang:2023vcv
$dl_{prop}=P_{\mu}u_{e}^{\mu}d\lambda=\frac{P_{t}}{g|P_{r}|}dr.$ (41)
Finally, by combining Eqs. (33) and (41), the total observed intensity can be calculated through the following expression Bambi:2013nla ; Wang:2023vcv
$I_{obs}=\int_{\gamma}\frac{g^{3}P_{t}}{r^{2}|P_{r}|}dr.$ (42)
In Fig. 13, we present the received intensities for both Schwarzschild and Schwarzschild-like BHs under different DM parameter configurations. The left panel of Fig. 13 displays the results with fixed $\rho_{s}=0.01$ for $r_{s}=0.1$, $0.15$, and $0.2$, while the right panel shows the cases with fixed $r_{s}=0.06$ for $\rho_{s}=0.1$, $0.15$, and $0.2$. Notably, the parameter-dependent variations in received intensity follow similar trends to those observed in static spherical accretion.
The corresponding optical appearances of BH surrounded by the infalling spherical accretion are shown in Fig. 14. Our analysis reveals that the shadow radius observed in the viewer’s frame consistently corresponds to the critical impact parameter for both static and radially infalling accretion scenarios. Moreover, the optical appearance of the infalling spherical accretion exhibits identical dependence on DM parameters as its static counterpart–both demonstrate progressively larger and darker central shadow regions with increasing DM parameter values. These results demonstrate that the optical appearance of BHs illuminated by different accretion models provides an effective observational diagnostic to distinguish Schwarzschild-like BHs from their classical Schwarzschild counterparts.
Figure 13: The variation of total observed intensity with impact parameter in the radially infalling spherical accretion. The left panel displays the received intensity for cases with fixed DM halo density $\rho_{s}=0.01$, where $r_{s}$ takes values of 0.1, 0.15, 0.2 and the Schwarzschild case, represented by blue, green, black and red curves respectively. The right panel shows the corresponding intensity comparison for fixed $r_{s}=0.06$ with $\rho_{s}$ values of 0.1, 0.15, 0.2 and the Schwarzschild case, similarly depicted using the same color scheme.
Figure 14: The BH images illuminated by the radially infalling spherical accretion. The left panel compares the optical appearances under the following configurations: (i) Schwarzschild BH (top left); (ii) Schwarzschild-like BHs with fixed $\rho_{s}=0.01$: $r_{s}=0.1$ (top right), $r_{s}=0.15$ (bottom right), and $r_{s}=0.2$ (bottom left). The right panel contrasts the optical appearances for: (i) Schwarzschild BH (top left); (ii) Schwarzschild-like BHs with fixed $r_{s}=0.06$: $\rho_{s}=0.1$ (top right), $\rho_{s}=0.15$ (bottom right), and $\rho_{s}=0.2$ (bottom left).
6 Conclusion